# BRIGHT: A globally distributed multimodal building damage assessment dataset with very-high-resolution for all-weather disaster response

### v1

@@ 1 Study areas and disaster events
We selected 12 disaster events across the globe for our BRIGHT dataset, as illustrated in Figure 2. Since both Capella and Umbra satellites were launched in 2020 and beyond, we focused on study areas where disasters occurred from 2020 onward. The selected regions are primarily in developing countries, where public administration and disaster response capacity tend to be weaker compared to developed nations, making international assistance more critical. The dataset covers five major types of natural disasters—earthquakes, storms (including hurricane and cyclone), wildfires, floods, and volcanic eruptions. Additionally, it includes man-made disasters such as explosions and armed conflicts.
@@
Ukraine | Armed conflict (AC) | 22 Mar. 2022 - 21 Sept. 2022 | 0.6 | Google Earth Capella | 848 | 56,770 |
@@
Myanmar | Cyclone (CC) | 14 May 2023 | 0.6 | Google Earth Capella | 126 | 8,052 |
@@
Acapulco, Mexico | Hurricane (HC) | 25 Oct. 2023 | 0.35-0.8 | Google Earth Capella | 325 | 18,437 |
Test events of IEEE DFC 2025 | N/A | N/A | N/A | N/A | N/A | N/A |
Total | - | - | - | - | >4,538 | >350,000 |
@@ 1.2 Explosion in Bata, 2021
On March 7, 2021, a series of four explosions occurred at the Cuartel Militar de Nkoantoma in Bata, Equatorial Guinea, caused by improperly stored explosives. The blasts led to at least 107 deaths, over 615 injuries, and widespread destruction throughout the city. A total of 243 structures were destroyed or severely damaged, displacing many residents. Around 150 families sought refuge in temporary shelters, while others stayed with relatives. Local hospitals treated more than 500 injured individuals, and the economic impact was severe, underscoring the dangers associated with improper storage of hazardous materials.
@@ 1.4 Earthquake in Haiti, 2021
On August 14, 2021, a magnitude 7.2 earthquake struck Haiti’s Tiburon Peninsula, primarily affecting the Nippes, Sud, and Grand’Anse departments. The disaster caused over 2,200 deaths, more than 12,200 injuries, and left thousands homeless. The economic losses were significant, estimated at over USD 1.5 billion. Approximately 137,500 buildings including homes, schools, and hospitals, were damaged or destroyed. As the deadliest natural disaster of 2021, the earthquake exacerbated Haiti’s existing challenges, including widespread poverty and political instability.
@@ 1.6 Armed Conflict in Ukraine, 2022
In February 2022, the conflict in Ukraine escalated significantly when Russian forces launched a full-scale invasion. The invasion has affected many parts of Ukraine, especially the eastern regions and the major cities, causing widespread destruction to homes, infrastructure, and industrial facilities. The conflict has resulted in thousands of civilian and military casualties and has caused the displacement of millions of people. Out of a population of 41 million, about 8 million Ukrainians were internally displaced, while more than 8.2 million fled the country by April 2023, creating Europe’s largest refugee crisis since World War II. The economic losses are substantial, with extensive damage to infrastructure, industrial output, and the broader economy.
@@ 1.10 Earthquake in Morocco, 2023
On September 8, 2023, a 6.9 magnitude earthquake struck Morocco’s Al Haouz Province, near Marrakesh, causing widespread devastation. The quake resulted in at least 2,960 deaths and 5,674 injuries, affecting over 2.8 million people, including 100,000 children. The earthquake damaged or destroyed 40,759 houses and 2,930 villages, with 19,095 additional houses collapsing. Significant destruction occurred in rural areas of the Atlas Mountains. Economic losses could reach up to 9% of Morocco’s GDP. Historic sites in Marrakesh, including parts of the Medina and several mosques, were severely damaged.
@@ 2.1 Construction of the dataset
The optical EO data in our dataset are mainly from Maxar’s Open Data program111https://www.maxar.com/open-data, while the SAR EO data are mainly from Capella Space and Umbra Space. Both Capella and Umbra data have two imaging modalities, i.e., Spotlight and Stripmap, respectively. The Spotlight mode has a higher spatial resolution but less coverage. In the region of interest, we preferred Spotlight mode if there is suitable data in the data provider’s inventory, otherwise, we chose Stripmap. The optical EO data consist of red, blue, and green bands, while the SAR EO data are mainly amplitude data in the VV or HH bands. For optical EO data, the digital number was converted to reflectance and then standardized to an 8-bit data format. For SAR imagery, after the data have been terrain-corrected, we used the data provider’s recommended method of pre-processing the amplitude data and finally converted the same to an 8-bit data format. Although both optical and SAR images are geocoded, there is still some pixel offset between them. Therefore, multiple experts manually aligned the paired optical and SAR data and cross-checked each other to ensure the precise registration between the two types of EO data.
@@
The labels in BRIGHT consist of two components: building polygons and post-disaster building damage attributes. Expert annotators manually labeled the building polygons, and all labels underwent independent visual inspections to ensure accuracy. Damage annotations were obtained from Copernicus Emergency Management Service222https://emergency.copernicus.eu and the United Nations Satellite Centre (UNOSAT)333https://unosat.org/products Emergency Mapping Products. These annotations were derived through visual interpretation of high-resolution optical imagery captured before and after the disasters by EO experts, supplemented by partial field visits. Since the number of damage levels varied across different events, we adopted the approach in (Adriano et al., 2021) to standardize them into three categories: intact, damaged, and destroyed. The damage annotations were provided as vector point files. By overlaying these points with the building polygons and assigning corresponding damage attributes, we generated the final building damage labels. To prevent geographic misallocations due to coordinate offsets, we unified the coordinate systems of the points and polygons, with a visual inspection performed prior to the final allocation.
@@
Figure 3 presents thumbnails of selected local areas from the 12 disaster events included in the BRIGHT dataset.
@@ 2.2 Statistics for BRIGHT
The basic information about BRIGHT, including disaster events, EO data, the number of corresponding EO tiles, and the total number of building pixels, is summarized in Table 2. After cropping the EO data into 10241024-pixel tiles, BRIGHT contains 4,538 multimodal image pairs (excluding events related to IEEE GRSS DFC 2025). As BRIGHT covers different disaster events located in different regions across the globe, this ensures the geographical diversity of its sample.
Key statistics of BRIGHT are illustrated in Figure 4. Figure 4a shows the pixel value distribution for optical and SAR images from one man-made disaster and three natural disasters. The varying geographical landscapes and land cover across different regions result in distinct means and standard deviations of pixel values, highlighting BRIGHT’s geographical diversity and making it a robust dataset for studying building damage assessment in diverse environments. To ensure trained deep learning (DL) models can accurately detect buildings and assess damage levels, it is crucial that the dataset includes a wide variety of building styles from different regions. Figure 4b shows that BRIGHT covers buildings at multiple scales, exhibiting a “long-tail” distribution. This multi-scale representation challenges DL models to develop the ability to capture features at varying scales, enhancing robustness and accuracy.
@@
Descriptor | RMSE (pixels) |
RIFT (Li et al., 2020) | 1.935 |
LNIFT (Li et al., 2022) | 0.905 |
SRIF (Li et al., 2023a) | 1.790 |
Average | 1.543 |
We also evaluated the alignment errors between optical and SAR EO data in BRIGHT. Table 3 presents the mean co-registration errors, measured in root mean square error (RMSE) of pixel displacement, obtained using three different multimodal image registration methods: RIFT (Li et al., 2020), SRIF (Li et al., 2023a), and LNIFT (Li et al., 2022). The evaluation resulted in an RMSE of about 1.5 pixels between optical and SAR EO data. This enhanced alignment provides a fundamental guarantee for accurate building damage assessment in BRIGHT.
@@ 2.3 Dataset division strategy
To train ML/DL models using BRIGHT and evaluate their generalizability, it is necessary to divide the dataset into a training set, validation set, and test set. Gerard et al. (2024) suggested that dividing the dataset on an event-by-event basis, rather than randomly across the entire dataset, provides a more accurate reflection of a model’s generalizability. Therefore, for the first 12 events listed in Table 2, we divide the corresponding data for each event into training, validation, and test sets in a ratio of 7:1:2. The sets obtained for each event are then merged to create the final training, validation, and test sets. In our experiments, DL models are trained using the training set, and the optimal hyperparameters (e.g., learning rate) are selected based on performance on the validation set. The generalization capability of the models is subsequently evaluated on the test set.
@@ 3.1 Problem statement and two solutions
One is to directly treat the building damage assessment task as a single semantic segmentation task (Adriano et al., 2021; Gupta and Shah, 2021). The pre- and post-event images are taken as inputs of the model, and then the final damage proxy map is directly predicted, which can be formalized as , where is is the pre-event imagery, is the post-event imagery, is a semantic segmentation model, is the obtained damage proxy map. In the context of this paper, is VHR optical imagery and is VHR SAR imagery.
The second solution adopts the task decoupling approach (Gupta et al., 2019; Zheng et al., 2021), which breaks down building damage assessment into two subtasks: the building localization task, i.e., separating the building from the background, and the damage classification task, i.e., focusing on the classification between different levels of damage. This solution can be formulated as and , where is the building localization map, is the damage classification map, and are models for building localization and damage classification tasks, respectively. and can be two separate models (Gupta et al., 2019) or a unified multi-task learning model (Zheng et al., 2021; Chen et al., 2022, 2024). The final building damage proxy map is obtained by combining the two outputs using a simple mask operation: .
@@
3.2 Deep learning architecture
We evaluate BRIGHT using several advanced deep network architectures from both the computer vision and remote sensing communities. Since the building damage assessment task can be considered a specialized semantic segmentation task, we selected two well-known segmentation networks from the computer vision field: UNet (Ronneberger et al., 2015) and DeepLabV3+ (Chen et al., 2018). Additionally, we included several state-of-the-art networks from the EO community: SiamAttnUNet (Adriano et al., 2021), SiamCRNN (Chen et al., 2020), ChangeOS (Zheng et al., 2021), DamageFormer (Chen et al., 2022), and ChangeMamba (Chen et al., 2024). These networks encompass a broad range of deep learning paradigms, including CNNs, RNNs, Transformers, and the more recent Mamba architecture.

### v2

@@ 1 Study areas and disaster events
We selected 14 disaster events across the globe for our BRIGHT dataset, as illustrated in Figure 2. Since both Capella and Umbra satellites were launched in 2020 and beyond, we focused on study areas where disasters occurred from 2020 onward. The selected regions are primarily in developing countries, where public administration and disaster response capacity tend to be weaker compared to developed nations, making international assistance more critical. The dataset covers five major types of natural disasters: earthquakes, storms (including hurricanes and cyclones), wildfires, floods, and volcanic eruptions. Additionally, it includes human-made disasters, such as accidental explosions and armed conflicts.
@@
| Boulder, USA | Wildfire (WF) | 30 Dec. 2021 - 01 Jan., 2022 | 0.6 | NAIP Capella | 77 | 8,365 |
| Ukraine | Armed conflict (AC) | 22 Mar. 2022 - 21 Sept. 2022 | 0.6 | Google Earth Capella | 513 | 56,770 |
@@
| Kyaukpyu, Myanmar | Cyclone (CC) | 14 May 2023 | 0.6 | Google Earth Capella | 126 | 8,052 |
@@
| Acapulco, Mexico | Hurricane (HC) | 25 Oct. 2023 | 0.35-0.8 | Google Earth Capella | 212 | 18,437 |
| Noto, Japan | Earthquake (EQ) | 01 Jan. 2024 | 0.5 | GSI Umbra | 79 | 8,153 |
| Total | - | - | 0.3-1 | - | 4,246 | 384,596 |
@@ 1.2 Explosion in Bata, 2021
On March 7, 2021, a series of four explosions occurred at the Cuartel Militar de Nkoantoma in Bata, Equatorial Guinea, caused by improperly stored explosives. The blasts led to at least 107 deaths, over 615 injuries, and widespread destruction throughout the city. A total of 243 structures were destroyed or severely damaged, displacing many residents. Around 150 families sought refuge in temporary shelters, while others stayed with relatives. Local hospitals treated more than 500 injured individuals, and the economic impact was severe, underscoring the dangers associated with the improper storage of hazardous materials.
@@ 1.4 Earthquake in Haiti, 2021
On August 14, 2021, a magnitude 7.2 earthquake struck Haiti’s Tiburon Peninsula, primarily affecting the Nippes, Sud, and Grand’Anse departments. The disaster caused over 2,200 deaths, more than 12,200 injuries, and left thousands homeless. The economic losses were significant, estimated at over USD 1.5 billion. Approximately 137,500 buildings, including homes, schools, and hospitals, were damaged or destroyed. As the deadliest natural disaster of 2021, the earthquake exacerbated Haiti’s existing challenges, including widespread poverty and political instability.
@@ 1.5 Volcano Eruption in La Palma, 2021
1.6 Wildfire in Colorado, 2021
The Marshall Fire, which started on December 30, 2021, in Boulder County, Colorado, became the most destructive wildfire in the state history in terms of destroyed buildings. Fueled by dry grass from an unusually warm and dry season, and winds up to 185 km/h, the fire killed two people and injured eight. It destroyed 1,084 structures, including homes, a hotel, and a shopping center, causing over US$2 billion in damage. More than 37,500 residents evacuated, and significant damage to public drinking water systems occurred.
@@ 1.7 Armed Conflict in Ukraine, 2022
In February 2022, the conflict in Ukraine escalated significantly when Russian forces launched a full-scale invasion. The invasion has affected many parts of Ukraine, especially the eastern regions and the major cities, causing widespread destruction to homes, infrastructure, and industrial facilities. The conflict has resulted in thousands of civilian and military casualties and has caused the displacement of millions of people. Out of a population of 41 million, about 8 million Ukrainians were internally displaced, while more than 8.2 million had fled the country by April 2023, creating Europe’s largest refugee crisis since World War II. The economic losses are substantial, with extensive damage to infrastructure, industrial output, and the broader economy.
@@ 1.11 Earthquake in Morocco, 2023
On September 8, 2023, a 6.9 magnitude earthquake struck Morocco’s Al Haouz Province near Marrakesh, causing widespread devastation. The quake resulted in at least 2,960 deaths and 5,674 injuries, affecting over 2.8 million people, including 100,000 children. The earthquake damaged or destroyed 40,759 houses and 2,930 villages, with 19,095 additional houses collapsing. Significant destruction occurred in rural areas of the Atlas Mountains. Economic losses could reach up to 9% of Morocco’s GDP. Historic sites in Marrakesh, including parts of the Medina and several mosques, were severely damaged.
@@ 1.13 Hurricane in Mexico, 2023
1.14 Earthquake in Noto, 2024
On January 1, 2024, a magnitude 7.5 earthquake struck the Noto Peninsula in Ishikawa Prefecture, Japan, reaching a maximum JMA seismic intensity of Shindo 7. The earthquake caused widespread destruction, particularly in the towns of Suzu, Wajima, Noto, and Anamizu, and triggered a 6.58-meter tsunami along the Sea of Japan coast. The disaster resulted in 339 deaths and over 1,300 injuries, with 156,526 structures damaged across nine prefectures. It was the deadliest earthquake in Japan since the 2011 Tōhoku disaster. As of February 20, 12,929 people remained in 521 evacuation centers. The event prompted Japan’s first major tsunami warning since 2011. Economic losses are estimated to be substantial, with significant impacts on infrastructure and communities across the affected regions.
@@ 2.1 Construction of the dataset
Figure 3 shows the flowchart of developing BRIGHT. The optical EO data in our dataset are mainly from Maxar’s Open Data program333https://www.maxar.com/open-data, while the SAR EO data are from Capella Space444https://www.capellaspace.com/earth-observation/gallery and Umbra Space555https://umbra.space/open-data/. Both Capella and Umbra data have two imaging modalities, i.e., Spotlight and Stripmap, respectively. The Spotlight mode has a higher spatial resolution but less coverage. In the region of interest, we preferred Spotlight mode if there is suitable data in the data provider’s inventory. Otherwise, we chose Stripmap. The optical EO data consist of red, blue, and green bands, while the SAR EO data are amplitude data in the VV or HH bands. For optical EO data, the digital number was converted to reflectance and then standardized to an 8-bit data format. For SAR imagery, after the data had been terrain-corrected, we used the data provider’s recommended method of pre-processing the amplitude data and finally converted the data to an 8-bit data format. Although both optical and SAR images are geocoded, there are still pixel offsets between them. Therefore, multiple EO experts manually aligned the paired optical and SAR data and cross-checked their results to ensure the precise registration between the two types of EO data. Figure 11 shows the selected control points on three disaster scenes.
| Class | Definition |
| Background (0) | All non-building pixels |
| Intact (1) | No visible signs of structural damage, water intrusion, shingle displacement, or burn marks. |
| Damaged (2) | Partial structural damage to the building, such as missing roof members, visible cracks, or partial wall/roof collapse. Buildings may be partially burned, surrounded by water or mud, or affected by nearby volcanic flows. |
| Destroyed (3) | Completely collapsed, burned, partially/completely covered by water/mud or no longer present |
@@
The labels in BRIGHT consist of two components: building polygons and post-disaster building damage attributes. Expert annotators manually labeled the building polygons, and all labels underwent independent visual inspections to ensure accuracy. Damage annotations were obtained from Copernicus Emergency Management Service666https://emergency.copernicus.eu, the United Nations Satellite Centre (UNOSAT) Emergency Mapping Products777https://unosat.org/products, and The Federal Emergency Management Agency (FEMA)888https://www.fema.gov. These annotations were derived through visual interpretation of high-resolution optical imagery captured before and after the disasters by EO experts, supplemented by partial field visits. To further improve damage information reliability, we performed a secondary validation and correction of these annotations. EO experts carefully reviewed the damage points using multi-temporal imagery available on platforms such as Google Earth and manually refined the labels when inconsistencies or misclassifications were observed. Since the number of damage levels varied across different events, we refered to the framework in FEMA’s Damage Assessment Operations Manual, EMS-98, the BDD dataset (Adriano et al., 2021), and the xBD dataset (Gupta et al., 2019) to standardize them into three categories: intact (1), damaged (2), and destroyed (3). Table 3 summarizes these definitions. The damage annotations were provided as vector point files. By overlaying these points with the building polygons and assigning corresponding damage attributes, we generated the final building damage labels. To prevent geographic misallocations due to possible coordinate offsets, we unified the coordinate systems of the points and polygons, with a visual inspection performed prior to the final allocation.
@@
Figure 4 presents thumbnails of selected local areas from the 14 disaster events included in the BRIGHT dataset.
@@ 2.2 Statistics for BRIGHT
The basic information about BRIGHT, including disaster events, EO data, the number of corresponding EO tiles, and the total number of building pixels, is summarized in Table 2. After cropping the EO data into 10241024-pixel tiles, BRIGHT contains 4,246 multimodal image pairs.
The key statistics of BRIGHT are illustrated in Figure 5. Figure 5a shows the pixel value distribution for optical and SAR images from one human-made disaster and three natural disasters. The varying geographical landscapes and land cover across different regions result in distinct means and standard deviations of pixel values, highlighting BRIGHT’s geographical diversity and making it a robust dataset for studying building damage assessment in diverse environments. To ensure trained deep learning (DL) models can accurately detect buildings and assess damage levels, it is crucial that the dataset includes a wide variety of building styles from different regions. Figure 5b shows that BRIGHT covers buildings at multiple scales, exhibiting a “long-tail” distribution. This multi-scale representation challenges DL models to develop the ability to capture features at varying scales, enhancing robustness and accuracy.
@@
| Descriptor | RMSE (All) | RMSE (Building) |
| RIFT (Li et al., 2020) | 1.125 | 1.138 |
| LNIFT (Li et al., 2022) | 0.857 | 0.825 |
| SRIF (Li et al., 2023a) | 1.090 | 1.054 |
| Average | 1.024 | 1.006 |
We also evaluated the registration accuracy between optical and SAR EO data in BRIGHT. Due to the absence of real ground truth, we introduce a proxy method to estimate registration errors, leveraging existing multimodal image descriptors. This approach is detailed in Appendix B. Table 4 reports the mean registration errors, measured as the root mean square error (RMSE) of pixel displacements, obtained using three representative multimodal image registration methods: RIFT (Li et al., 2020), SRIF (Li et al., 2023a), and LNIFT (Li et al., 2022). The overall average RMSE is approximately 1.024 pixels, with a lower error of 1.006 pixels specifically within building regions. Accurate registration results ensure spatial consistency across modalities, providing a solid foundation for accurate building damage assessment in BRIGHT.
@@ 2.3 Dataset division strategy
To train ML/DL models using BRIGHT and evaluate their generalizability, it is necessary to divide the dataset into a training set, validation set, and test set. Gerard et al. (2024) suggested that dividing the dataset on an event-by-event basis, rather than randomly across the entire dataset, provides a more accurate reflection of a model’s generalizability. Therefore, for the 14 events listed in Table 2, we divide the corresponding data for each event into training, validation, and test sets in a ratio of 7:1:2. The sets obtained for each event are then merged to create the final training, validation, and test sets. In our experiments, DL models are trained using the training set, and the optimal hyperparameters (e.g., learning rate) are selected based on performance on the validation set. The generalization capability of the models is subsequently evaluated on the test set.
In addition to the above standard ML division, we also introduce a cross-event transfer setup to better evaluate the ability of models to generalize across disaster events, a critical challenge in real-world applications where models are expected to handle unseen disaster types and locations. This setup comprises two key divisions:
Zero-shot setup: This setting mimics a real-world scenario where a newly occurring disaster must be analyzed without any prior labeled data from the same event. We isolate one event as an unseen test set while using the remaining 13 events for training and validation. This setting evaluates the cross-event generalization ability of models, testing how well they can transfer learned knowledge from previous disasters to an entirely new event. Due to the high variability of disaster types and geographies, this setup is inherently challenging, as models trained on past disasters may struggle to assess damage patterns in a previously unseen event accurately.
One-shot setup: Recognizing the difficulty of the zero-shot setup, we introduce a one-shot setup, acknowledging that in real-world disaster response, a small number of labeled samples from the new disaster can be obtained. In this setting, a limited subset of labeled data (one image for training and one image for validation) from the target disaster event is incorporated into the training process. At the same time, the majority of the test set remains unseen. This setup evaluates the model’s ability to leverage a minimal amount of manually labeled data to improve disaster-specific adaptation.
@@ 3.1 Problem statement and two solutions
One is to directly treat the building damage assessment task as a single semantic segmentation task (Adriano et al., 2021; Gupta and Shah, 2021). The pre- and post-event images are taken as inputs of the model, and then the final damage map is directly predicted, which can be formalized as , where is is the pre-event imagery, is the post-event imagery, is a semantic segmentation model, is the obtained damage map. In the context of this paper, is VHR optical imagery and is VHR SAR imagery.
The second solution adopts the task decoupling approach (Gupta et al., 2019; Zheng et al., 2021), which breaks down building damage assessment into two subtasks: the building localization task, i.e., separating the building from the background, and the damage classification task, i.e., focusing on the classification between different levels of damage. This solution can be formulated as and , where is the building localization map, is the damage classification map, and are models for building localization and damage classification tasks, respectively. and can be two separate models (Gupta et al., 2019) or a unified multi-task learning model (Zheng et al., 2021; Chen et al., 2022a, 2024). The final building damage map is obtained by combining the two outputs using a simple mask operation: .
@@
3.2 Benchmark suites
We evaluate BRIGHT using several advanced deep network architectures from both the computer vision and remote sensing communities. Since the building damage assessment task can be considered a specialized semantic segmentation task, we selected two well-known segmentation networks from the computer vision field: UNet (Ronneberger et al., 2015) and DeepLabV3+ (Chen et al., 2018). Additionally, we included several state-of-the-art networks from the EO community: SiamAttnUNet (Adriano et al., 2021), SiamCRNN (Chen et al., 2020), ChangeOS (Zheng et al., 2021), DamageFormer (Chen et al., 2022a), and ChangeMamba (Chen et al., 2024). These networks encompass a broad range of representative deep learning paradigms, including CNNs, RNNs, Transformers, and the more recent Mamba architecture.
@@ 3.2 Benchmark suites
Beyond supervised deep learning, BRIGHT also enables the evaluation of broader strategies commonly explored in EO and vision communities:
Unsupervised domain adaptation (UDA) methods for the zero-shot transfer setup, enabling models to transfer knowledge across disaster events with no labeled samples from the target event.
Semi-supervised learning (SSL) approaches for the one-shot transfer setup, leveraging a small number of labeled samples and the remaining unlabeled samples from new disaster events to refine model adaptation.
Unsupervised multimodal change detection (UMCD) methods, which exploit the modality-independent relationship in optical and SAR data to detect land-cover changes without requiring manual annotations.
Unsupervised multimodal image matching (UMIM) methods, which aim to learn modality-independent features to enable automatic registration of multimodal data without relying on manual alignment.
| Method | () | () | Final OA () | Final mIoU () | IoU per class () | |||
| Background | Intact | Damaged | Destroyed | |||||
| UNet | 87.97 | 72.24 | 95.47 | 64.94 | 96.19 | 71.27 | 39.13 | 53.17 |
| DeepLabV3+ | 87.00 | 70.33 | 95.43 | 64.80 | 95.98 | 70.98 | 37.53 | 54.69 |
| SiamAttnUNet | 88.16 | 70.13 | 95.45 | 64.26 | 96.14 | 71.90 | 35.06 | 53.92 |
| SiamCRNN | 89.45 | 72.02 | 95.76 | 65.73 | 96.48 | 73.44 | 38.91 | 54.18 |
| ChangeOS | 89.60 | 71.88 | 95.84 | 65.98 | 96.54 | 73.85 | 38.99 | 54.53 |
| DamageFormer | 90.29 | 72.51 | 96.13 | 67.09 | 96.87 | 75.04 | 39.86 | 56.59 |
| ChangeMamba | 90.90 | 72.70 | 96.22 | 67.63 | 96.96 | 75.59 | 40.05 | 57.91 |